OpenStudy (anonymous):

Solve: Dx [(cotxcscx)^2]

OpenStudy (mathmale):

You may find this problem easier to do if you'd first re-write the cot x and csc functions in terms of sin x and cos x:\[\frac{ d }{ dx }(\frac{ \cos x }{ \sin x }*\frac{ 1 }{ \sin x })^2\]

OpenStudy (anonymous):

what's next?